# Improving Representation of High-frequency Components for Medical Foundation Models

### v2

@@ 4.5 Ablation study
We gradually perform the ablation study of the components in Frepa to evaluate their respective contributions. The validation is performed on datasets including Pneumonia MNIST [12] for classification, Lung cancer [37] for lesion detection, and OCTA arteries and veins [48] for segmentation. The four most critical components of Frepa examined are frequency dual-component masking (FM), equalized-histogram image-domain masking (IM), HFL loss, and adversarial training (AT). For the ablation, these components are individually replaced or removed as follows: FM branch is deleted, IM is replaced by zero-patch masking (Fig. 3, second column), HFL loss, and AT are excluded during the pretraining phase.
| Pneumonia MNIST | Lung cancer | OCTA artery | OCTA vein | |||
| Metric | ACC | AUC | Detetcion | IoU | DSC() | DSC() |
| Frepa+ViT | 0.934 | 0.970 | 0.951 | 0.656 | 78.00 | 76.15 |
| w/o FM | 0.921 | 0.965 | 0.916 | 0.590 | 73.48 | 71.71 |
| w/o IM | 0.931 | 0.972 | 0.919 | 0.607 | 76.31 | 73.89 |
| w/o HFL | 0.927 | 0.967 | 0.939 | 0.624 | 74.69 | 73.11 |
| w/o AT | 0.933 | 0.970 | 0.945 | 0.632 | 77.70 | 76.03 |
| Frepa+SwinT | 0.932 | 0.981 | 0.948 | 0.677 | 79.08 | 77.35 |
| w/o FM | 0.913 | 0.971 | 0.905 | 0.604 | 73.96 | 72.40 |
| w/o IM | 0.926 | 0.977 | 0.919 | 0.621 | 76.58 | 76.65 |
| w/o HFL | 0.930 | 0.978 | 0.928 | 0.624 | 76.23 | 75.90 |
| w/o AT | 0.932 | 0.980 | 0.933 | 0.655 | 78.82 | 76.28 |
The results of the ablation study are presented in Table V. Removing each component leads to a decline in model performance across all tasks. Notably, the FM has the most significant impact; its removal causes the performance to degrade to be comparable to that of the MAE, suggesting a failure in high-frequency representation. Similarly, the results indicate that the zero-patch masking is less effective than our proposed equalized-histogram IM strategy. Additionally, the exclusion of HFL loss and AT also contributes to a reduction in performance. These findings are consistent for both ViT and SwinT architectures. In summary, the masking strategy and loss function designs are crucial components that significantly enhance the performance of Frepa.
@@ 4.6 Frequency representation in image embedding
MSE loss is employed during training. Subsequently, we compare the reconstructed images with the raw images across different frequency components, utilizing the RMSE as the metric for reconstructed ratios.
where represents low-pass, medium-pass, and high-pass filtering, respectively. We compared Frepa, implemented with both ViT and SwinT, against two other autoencoder methods: MAE and MFM. The results of this comparison are presented in Fig. 6b. It is evident that Frepa consistently outperforms the other autoencoder methods in medium-frequency and high-frequency components while maintaining comparable performance in low-frequency components. This underscores Frepa’s superior ability to represent and preserve high-frequency information without compromising low-frequency representation. These findings align with our experiments on downstream tasks, providing a direct and intuitive evaluation of the level of embedding spaces.

### v3

@@ 4.5 Ablation study
We perform the ablation study of the components in Frepa to evaluate their respective contributions. The validation is performed on datasets including Pneumonia MNIST [12] for classification, Lung tumor [74] for lesion detection, and OCTA arteries and veins [84] for segmentation. The seven most critical components of Frepa examined are high-frequency masking (HM), low-frequency perturbation (LP), frequency dual-component masking (FM), equalized-histogram image-domain masking (IM), HFL loss, embedding consistency loss (ConL), and gradient loss (GradL). For the ablation, these components are individually replaced or removed as follows: HM and LP are separately performed, FM branch is deleted, IM is replaced by zero-patch masking (Fig. 3, second column), HFL, ConL, and GradL are excluded during the pretraining phase.
Frepa+ViT | w/o HM | w/o LP | w/o FM | w/o IM | w/o HFL | w/o ConL | w/o GradL | ||
ACC | 0.934 | 0.931 | 0.927 | 0.921 | 0.931 | 0.927 | 0.933 | 0.933 | |
Pneumonia MNIST | AUC | 0.970 | 0.968 | 0.966 | 0.965 | 0.972 | 0.967 | 0.970 | 0.968 |
Detection | 0.951 | 0.943 | 0.938 | 0.916 | 0.919 | 0.939 | 0.945 | 0.949 | |
Lung tumor | IoU | 0.656 | 0.634 | 0.623 | 0.590 | 0.607 | 0.624 | 0.632 | 0.655 |
OCTA artery | DSC() | 78.00 | 73.61 | 76.48 | 73.48 | 76.31 | 74.69 | 77.70 | 74.95 |
OCTA vein | DSC() | 76.15 | 73.04 | 73.55 | 71.71 | 73.89 | 73.11 | 76.03 | 74.29 |
Frepa+SwinT | w/o HM | w/o LP | w/o FM | w/o IM | w/o HFL | w/o ConL | w/o GradL | ||
ACC | 0.932 | 0.929 | 0.921 | 0.913 | 0.926 | 0.930 | 0.932 | 0.932 | |
Pneumonia MNIST | AUC | 0.981 | 0.980 | 0.972 | 0.971 | 0.977 | 0.978 | 0.980 | 0.979 |
Detection | 0.948 | 0.942 | 0.935 | 0.905 | 0.919 | 0.928 | 0.933 | 0.941 | |
Lung tumor | IoU | 0.677 | 0.658 | 0.642 | 0.604 | 0.621 | 0.624 | 0.655 | 0.676 |
OCTA artery | DSC() | 79.08 | 75.08 | 77.73 | 73.96 | 76.58 | 76.23 | 78.82 | 76.70 |
OCTA vein | DSC() | 77.35 | 73.11 | 77.17 | 72.40 | 76.65 | 75.90 | 76.28 | 76.09 |
Frepa+ConvNeXt | w/o HM | w/o LP | w/o FM | w/o IM | w/o HFL | w/o ConL | w/o GradL | ||
ACC | 0.934 | 0.931 | 0.927 | 0.921 | 0.931 | 0.927 | 0.933 | 0.933 | |
Pneumonia MNIS | AUC | 0.970 | 0.968 | 0.966 | 0.965 | 0.972 | 0.967 | 0.970 | 0.970 |
Detection | 0.939 | 0.933 | 0.925 | 0.918 | 0.892 | 0.905 | 0.914 | 0.927 | |
Lung tumor | IoU | 0.657 | 0.651 | 0.629 | 0.625 | 0.593 | 0.599 | 0.608 | 0.653 |
OCTA artery | DSC() | 79.80 | 76.19 | 78.45 | 75.13 | 78.12 | 76.75 | 78.53 | 76.97 |
OCTA vein | DSC() | 77.00 | 72.91 | 75.39 | 72.50 | 75.48 | 74.00 | 76.11 | 75.75 |
The results of the ablation study are presented in Table IX. Removing each component leads to a decline in model performance across all tasks. Notably, the FM has the most significant impact; its removal causes the performance to degrade to be comparable to that of the MAE, suggesting a failure in high-frequency representation. Within the FM strategy, high-frequency masking and low-frequency perturbation serve distinct functions: for tasks that primarily depend on frequency features, such as Pneumonia MNIST classification, low-frequency perturbation plays a more critical role, whereas for tasks requiring fine-grained details, such as OCTA vessel segmentation, high-frequency masking becomes more essential. The results also indicate that the zero-patch masking is less effective than our proposed equalized-histogram IM strategy. Additionally, the exclusion of HFL loss, consistency loss, and gradient loss also contributes to a general reduction in performance. These findings are consistent for ViT, SwinT, and ConvNeXt architectures. In summary, the masking strategy and loss function designs are crucial components that significantly enhance the performance of Frepa.
@@ 4.6 Frequency representation in image embedding
MSE loss is employed during training. Subsequently, we compare the reconstructed images with the raw images across different frequency components, utilizing 1-RMSE as the metric to measure the reconstruction similarity.
where represents low-pass, medium-pass, and high-pass filtering, respectively. A higher value of indicates better restoration. We compared Frepa, implemented with ViT, SwinT, and ConvNeXt, against two other autoencoder methods: MAE and MFM. The results of this comparison are presented in Fig. 7b. It is evident that Frepa consistently outperforms the other autoencoder methods in medium-frequency and high-frequency components while maintaining comparable performance in low-frequency components. This underscores Frepa’s superior ability to represent and preserve high-frequency information without compromising low-frequency representation. Among the three backbones, the performance of Swin Transformer and ConvNeXt shows slightly better preservation compared to ViT under medium and high-pass filters. This aligns with the observation that the pure ViT architecture tends to exhibit weaker representation capabilities in high-frequency domains. These findings align with our experiments on downstream tasks, providing a direct and intuitive evaluation of the level of embedding spaces.